OpenStudy (anonymous):

Jennifer wants to close off a rectangular area in her lawn so that she can leave her dog outside. She buys enough fencing to close off 252 square feet of lawn. If the length of the enclosure is 21 feet, what is the width? 1. 231 feet 2. 12 feet 3. 14 feet 4. 273 feet

OpenStudy (anonymous):

What's the formula of rectangular area?

OpenStudy (anonymous):

w x h

OpenStudy (anonymous):

A = w * L Do you have area and length?

OpenStudy (anonymous):

252 x 21

OpenStudy (anonymous):

A = w * L 252 = 21 w Can you find width now?

OpenStudy (anonymous):

12?

OpenStudy (anonymous):

Can you test your result?

OpenStudy (anonymous):

21 x 12 = 252

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

Is it clear to you?

OpenStudy (anonymous):

yep, thanks a lot! xx